A mass of 7gm draws up mass of 5gm connected to it by string passing over smooth fixed pulley.At the end of first second the string is cut.find the velocity of mass 7gn at the end of next second

Grade:12th pass

1 Answers

harshin nechooli
5 years ago
Here initial is the blocks connected by the string then their.  a. = 5÷3      v= at =5÷3 
then.   next 1 second. g acts on block  
      v= u. +at    = 35÷3